OneChicago has just introduced a single stock futures contract on Brandex stock, a company that currenty pays no dividends. Eacth contract calls for delivery of 1,400 shares of stock in 1 year. The T-bill rate is 5% per year a. If Brandex stock now sells at $150 per share, what should the futures price be? (Round your answer to 2 decimal places.) Futures price b. If the Brandex price drops by 3%, what will be the new futures price and the change in the investors margin account? (Round Futures price (new) answer to 3 decimal places and other answer to the nearest dollar amount. Negative amount should be indicated by a minus sign.) Futures price (new) Change in the investors margin account
media%2F1a1%2F1a16448a-202b-4f58-ab1d-71
Answer #1
A Future Contract is an agreement that gives the investor right to buy or sell the underlying assets at a fixed price at a future date.
Single stock future (SSF'S) are future contacts on individually listed shares. They are standardized Contracts with set specification regarding size, expiry dates and delivery.
The future price (F0) IS CALCULATED USING THE FOLLOWING FORMULA :
F0= S0(1+RF-D)^T      …….(1)
HERE, S0 IS CURRENT PRICE OF STOCK, T IS TE MATURITY PERIOD , RF IS THE RATE OF INTEREST AND D IS THE DIVIDEND YIELD.
OneChicago has just introduced a single-stock futures contract on Brandex stock, a company that currently pays no dividends. Each contract calls for delivery of 1,400 shares of stock in 1 year. The T-bill rate is 5% per year.
a. If Brandex stock now sells at $150 per share, what should the futures price be?
FROM (1)
F0= S0(1+RF-D)^T
F0= 150 *(1+ .05 -0)^1
F0= 157.5
b. If the Brandex price drops by 3%, what will be the change in the futures price and the change in the investor’s margin account?
New Spot = CURRENT PRICE OF THE STOCK (1 – PRICE DROPS )
New Spot = $150 (1 – 0.03)
New Spot = $150 x 0.97
New Spot = $145.50
New Futures = S0(1+RF-D)^T
New Futures = $145.50 (1.05)
New Futures = $152.775
The long investor loses
= future price befor price drops - price drops after prize drops
= $157.05-$152.775
= $4.725 per share
or 6615 per contract { $4.725(1400)}
c. If the margin on the contract is $13,200, what is the percentage return on the investor’s position?
Percentage return
= investor profit or loss from the contract / margin on the contract
= 6615 / 13200
= 50.11%